# A Concept-Centric Approach to Multi-Modality Learning

### v1

@@ 3.2 Learning Projection Models
Defining Projection Models. Decoupled from the abstract concept space, each modality-specific projection model can be viewed as a mapping function $f_{*}:\mathcal{X_{*}}\rightarrow\mathcal{K}$ that generates a box representation in $\mathcal{K}$ for each input from its modality-specific sample space $\mathcal{X_{*}}$ of an unspecified modality denoted by *. This projection onto $\mathcal{K}$ allows interactions between specific objects from $\mathcal{X_{*}}$ and abstract concepts in $\mathcal{C}$. Specifically, given a modality-specific input $x^{*}_{i}\in\mathcal{X_{*}}$, its representation in $\mathcal{K}$ can be obtained by $f_{*}(x^{*}_{i};\theta)=\Omega_{i}^{*}$ where $\Omega_{i}^{*}\subset\mathcal{K}$ follows the same definition of $\Omega_{y}\in\mathcal{C}$. With this representation made available, the probability that an object is associated with a concept $c$ can be naturally described by an entailment probability of $P(y|x_{i}^{*})=P(\Omega_{y}|\Omega_{i}^{*})$.
Adapting to Concept Space. Given $f_{*}$’s corresponding modality training set $\mathcal{D}_{*}$, not only should the projection produced for an input $x^{*}_{i}$ entail a single concept $y$, but it should also entail all other concepts related to $x^{*}_{i}$. In other words, the projection $\Omega^{*}_{i}$ for $x^{*}_{i}$ should lie at the intersection of a set of concepts that can describe $x^{*}_{i}$. Namely, the most optimal projection for $x^{*}_{i}$ should maximize the entailment probability of $P(\bigcap_{y_{j}\in\bm{y}_{i}}y_{j}|x^{*}_{i})$.
To drive projection models to produce this most optimal projection, we use a combination of a binary cross-entropy loss on attribute concepts $\mathcal{Y}^{attr}\subset\mathcal{Y}$:
$\begin{split}\ell_{attr}(\bm{y},\Omega_{*})&=\frac{1}{|\mathcal{Y}^{attr}|}\sum_{y\in\mathcal{Y}^{attr}}\mathbb{I}(y\in\bm{y})[-w\cdot\log P(\Omega_{y}|\Omega_{*})]\\ &+\mathbb{I}(y\notin\bm{y})[\log(1-P(\Omega_{y}|\Omega_{*})]\end{split}$ (1)
@@
and a multi-class cross-entropy loss with SoftMax on category concepts $\mathcal{Y}^{cat}\subset\mathcal{Y}$:
$\ell_{cat}(\bm{y},\Omega_{*})=-\log\frac{\exp{P(\Omega_{y^{cat}}|\Omega_{*})}}{\sum_{y^{\prime}\in\mathcal{Y}^{cat}}\exp{P(\Omega_{y^{\prime}}|\Omega_{*})}}$ (2)
(where $y^{cat}\in\bm{y}$)
@@
$\begin{split}\mathcal{L}_{A}(\theta_{A};\mathcal{D}_{A})=\frac{1}{|\mathcal{D}_{A}|}\sum_{(x,\bm{y})\in\mathcal{D}_{A}}\ell_{attr}(\bm{y},f_{A}(x;\theta_{A}))+\ell_{cat}(\bm{y},f_{A}(x;\theta_{A}))\end{split}$ (3)
@@
3.3 Cross Modality Joint Training
To allow probabilistic analysis for cross-modality tasks, we introduce a joint training stage that encourages different projection models to produce projections that overlap with each other’s for the same object. This joint training stage is lightweight since modality-specific projection models have already been trained and adapted to a unified concept space. It requires very modest resources, with convergence occurring within a few hundred training steps, as indicated in Fig. 5 of Appendix. Subsequently, this design with demonstrated efficiency allows the effortless incorporation of new projection models into our proposed framework, mirroring humans’ ability to learn and link knowledge across modalities in a fast and efficient manner. Specifically, consider a system with two modalities, A and B, as an example. The training dataset would be denoted as $\mathcal{D}_{A\cup B}=\{(x^{A}_{i},x^{B}_{i},\bm{y}_{i})\}_{i=1}^{N}$, and the training objective for this joint training stage is defined as:
($\begin{split}\mathcal{L}&{}_{\text{joint}}(\theta_{A},\theta_{B};\mathcal{D}_{A\cup B})=\frac{1}{2|\mathcal{D}_{A\cup B}|}\sum_{(x_{A},x_{B},\bm{y})\in\mathcal{D}_{A\cup B}}\\ &P(f_{A}(x_{A};\theta_{A})|f_{B}(x_{B};\theta_{B}))+P(f_{B}(x_{B};\theta_{B})|f_{A}(x_{A};\theta_{A}))\end{split}$ (4)
The overall training objective becomes a combination of modality-specific projection losses and this joint training loss. Optionally, optimization can also include parameters from $\mathcal{C}$, so that the abstract knowledge learned in the concept space is adjusted based on modality-specific information. Then the objective becomes $L_{\text{joint}}^{\prime}=L_{\text{joint}}+\beta L_{\mathcal{C}}$ where $L_{\mathcal{C}}$ denotes the KL divergence loss of the concept space.
@@ 4 Implementation and Experiments
We base our evaluation on three datasets: CLEVR (Johnson et al., 2017a), COCO (Lin et al., 2014), and GQA (Hudson & Manning, 2019) where their concepts are formed from original and supplemental annotations. Both attribute and categorical concepts are present in COCO and GQA whereas CLEVR only contains attribute concepts. More details on the datasets and preprocessing steps can be found in Appendix B. Our experiments follow the same train and validation splits as the original datasets. The proposed framework is pretrained on the train sets and tested on the validation sets.
@@ 4.1 Pretraining
Concept Space. To ensure that each concept box always has a valid set of lower boundaries smaller than its upper boundaries, we use two vectors, $(\omega_{min,y},\omega_{\Delta,y})=\Omega_{y}$, instead of $(\omega_{min},\omega_{max})$ to represent a box in our actual experiments, where $\omega_{\Delta}\in\mathcal{K}_{\geq 0}$ is restricted to non-negative values. A box’s upper boundaries can be obtained by $\omega_{max}=\omega_{min}+\omega_{\Delta}$. We set the dimension of $\mathcal{K}$ to 50, based on empirical experiments. Initial values for $\mathcal{C}$ are sampled from two uniform distributions. As for the negative sampling method, in CLEVR, the only negative concept pairs come from combinations of concepts residing in the same-attribute families, such as (red, blue). For COCO and GQA, negative samples are randomly selected from all concepts. The concept space is trained for two epochs for each dataset with a batch size of 256 using an AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $10^{-3}$. The training of this concept space can be completed quickly as there are only thousands of parameters for a moderately-sized concept space.
Projection Models. In adapting our framework to the datasets featuring vision and natural language modalities, we incorporate a vision projection model $f_{\text{vision}}$ based on a Vision Transformer encoder (Dosovitskiy et al., 2020) and a natural language projection model $f_{\text{NL}}$ based on a BERT encoder (Devlin et al., 2018). Both models utilize their encoders’ outputs on [CLS] tokens to generate projection boxes in $\mathcal{K}$. The outputs $e$ with a dimension of 768 are divided into two equal chunks, $h_{min}$ and $h_{\Delta}$, each with a dimension of 384. These chunks are then input into two fully connected layers to produce $\omega_{min}$ and $\omega_{\Delta}$ for their respective projection boxes. To ensure $\omega_{\Delta}$ is always a non-negative vector, an additional ReLU layer is applied. The complete projection process for inputs from the vision modality is outlined in Algorithm 1.
@@
0:  $\omega^{\text{vision}}_{\Delta}\in\mathcal{K}_{\geq 0},\Omega^{\text{vision}}\subset\mathcal{K}$
@@
  $h_{min}^{\text{vision}},h_{\Delta}^{\text{vision}}\leftarrow\text{split}(e_{\text{vision}})$
  $\omega^{\text{vision}}_{min}\leftarrow\text{Linear}_{min}^{\text{vision}}(h_{min})$
@@
  $\Omega^{\text{vision}}=(\omega^{\text{vision}}_{min},\omega^{\text{vision}}_{\Delta})$
@@
For each object $i$ in the CLEVR dataset, its attribute prediction for a specific attribute family $\bm{z}$ (e.g., color) is generated by $\bar{y}_{i}^{z}=\text{argmax}_{y\in\bm{z}}P(\Omega_{y}|\Omega_{i})$. For each object $i$ in COCO and GQA, a threshold is applied to $P(\Omega_{y}|\Omega_{i}),y\in\mathcal{Y}^{attr}$ to obtain attribute predictions, and category prediction is generated by $\bar{y}_{i}^{\text{cat}}=\text{argmax}_{y\in\mathcal{Y}^{cat}}P(\Omega_{y}|\Omega_{i})$.
@@
We establish a baseline by replacing the concept space with a traditional Multilayer Perceptron (MLP) at the classification head of $f_{vision}$. Additionally, we implement the vision-modality projection model using a ResNet model (He et al., 2015) as the backbone to showcase the flexibility of the proposed framework. Results summarized in Table 1 show that our proposed framework achieves comparable performance to traditional models while leveraging a novel concept space with interpretable learned knowledge.
Backbone Method CLEVR COCO GQA
Accuracy Accuracy F1 Score Accuracy F1 Score
ResNet Baseline $0.997_{\pm 1.0e^{-4}}$ $0.897_{\pm 1.7e^{-3}}$ $0.625_{\pm 2.4e^{-3}}$ $0.733_{\pm 4.5e^{-3}}$ $0.401_{\pm 2.8e^{-3}}$
$f_{vision}$ $0.990_{\pm 2.7e^{-3}}$ $0.900_{\pm 7.8e^{-4}}$ $0.621_{\pm 2.1e^{-3}}$ $0.724_{\pm 1.7e^{-3}}$ $0.429_{\pm 3.1e^{-3}}$
ViT Baseline $0.999_{\pm 3.9e^{-5}}$ $0.956_{\pm 2.0e^{-3}}$ $0.663_{\pm 1.7e^{-3}}$ $0.841_{\pm 2.3e^{-3}}$ $0.567_{\pm 1.7e^{-3}}$
$f_{vision}$ $0.999_{\pm 4.0e^{-5}}$ $0.955_{\pm 1.4e^{-3}}$ $0.658_{\pm 2.4e^{-3}}$ $0.839_{\pm 3.0e^{-3}}$ $0.574_{\pm 1.1e^{-3}}$
Table 1: A comparison with baseline models on classification performance of vision-modality inputs. Category concepts are evaluated with accuracy (%) and attribute concepts with f1 score. 2-sigma errors over five trails of experiments are reported
@@
Apart from featuring a concept-centric learning scheme, the proposed framework can also learn modality-specific knowledge faster by referencing learned knowledge from the modality-agnostic concept space as indicated in Fig. 2. This more natural learning process of our framework bridges the efficiency gap between traditional machine learning methods, which often demand extensive data, and human learning, which excels at adeptly and efficiently extracting modality-specific representations and associating them with mental entities of abstract knowledge. To fully evaluate the impact of this transparent, modality-agnostic concept space on the learning of modality-specific projection models, we conduct an ablation study on it in Sec. 5.
@@
Figure 2: Learning curves of proposed projection models and baseline models. Shaded area in plots represents 2-sigma errors over five trails of experiments. During the learning process, the proposed vision-modality projection model converges faster compared to the baseline thanks to the universal concept space that already has abstract knowledge embedded in it. This faster learning process of our framework bridges the efficiency gap between traditional machine learning methods, which require a huge amount of data, and human learning that excels at extracting modality-specific representations and linking them to mental entities of abstract knowledge.
@@
Projection models for the natural-language modality achieve highly accurate performance ($\geq 99\%$) thanks to the clearly structured description sentences. Further implementation and training details of projection models can be found in Appendix C.

### v2

@@ 3.2 Learning Projection Models
Defining Projection Models. Decoupled from the abstract concept space, each modality-specific projection model can be viewed as a mapping function $f_{*}\mathrel{\mathop{\ordinarycolon}}\mathcal{X_{*}}\rightarrow\mathcal{K}$ that generates a box representation in $\mathcal{K}$ for each input from its modality-specific sample space $\mathcal{X_{*}}$ of an unspecified modality denoted by *. This projection onto $\mathcal{K}$ allows interactions between specific objects from $\mathcal{X_{*}}$ and abstract concepts in $\mathcal{C}$. Specifically, given a modality-specific input $x^{*}_{i}\in\mathcal{X_{*}}$, its representation in $\mathcal{K}$ can be obtained by $f_{*}(x^{*}_{i};\theta)=\Omega_{i}^{*}$ where $\Omega_{i}^{*}\subset\mathcal{K}$ follows the same definition of $\Omega_{y}\subset\mathcal{C}$. With this representation made available, the probability that an object is associated with a concept $y$ can be naturally described by an entailment probability of $P(y\mid x_{i}^{*})=P(\Omega_{y}\mid\Omega_{i}^{*})$.
Adapting to the Concept Space. Given the training set $\mathcal{D}_{*}$ corresponding to a modality marked by *, the projection produced for an input $x^{*}_{i}$ should entail not only a single concept $y$ but also all other concepts associated with $x^{*}_{i}$. In other words, the projection $\Omega^{*}_{i}$ for $x^{*}_{i}$ should lie at the intersection of the set of concepts describing $x^{*}_{i}$. Thus, the optimal projection for $x^{*}_{i}$ should maximize the entailment probability $P(\bigcap_{y_{j}\in\bm{y}_{i}}y_{j}\mid x^{*}_{i})$.
In our experimental datasets, concepts fall into two types: attribute concepts and category concepts. A sample may possess multiple positive attributes but is associated with exactly one mutually exclusive category. As a result, attribute concepts require independent binary classification, whereas category concepts require a single multi-class decision. Naturally, a binary cross-entropy loss is applied to attribute concepts, and a SoftMax cross-entropy loss is applied to category concepts. Further details about the datasets and the nature of these concept types are provided in Sec. 4.
To drive projection models to produce the optimal projection described above, we use a combination of a binary cross-entropy loss on attribute concepts $\mathcal{Y}^{\text{attr}}\subset\mathcal{Y}$:
$\begin{split}\ell_{\text{attr}}(\bm{y},\Omega_{*})&=\frac{1}{|\mathcal{Y}^{\text{attr}}|}\sum_{y\in\mathcal{Y}^{\text{attr}}}\mathbb{I}(y\in\bm{y})[-w\cdot\log P(\Omega_{y}\mid\Omega_{*})]\\ &+\mathbb{I}(y\notin\bm{y})[\log(1-P(\Omega_{y}\mid\Omega_{*})]\end{split}$ (1)
@@
and a multi-class cross-entropy loss with SoftMax on category concepts $\mathcal{Y}^{\text{cat}}\subset\mathcal{Y}$:
$\ell_{\text{cat}}(\bm{y},\Omega_{*})=-\log\frac{\exp{P(\Omega_{y^{\text{cat}}}\mid\Omega_{*})}}{\sum_{y^{\prime}\in\mathcal{Y}^{\text{cat}}}\exp{P(\Omega_{y^{\prime}}\mid\Omega_{*})}}$ (2)
(where $y^{\text{cat}}\in\bm{y}$)
@@
$\begin{split}\mathcal{L}_{A}(\theta_{A};\mathcal{D}_{A})=\frac{1}{|\mathcal{D}_{A}|}\sum_{(x,\bm{y})\in\mathcal{D}_{A}}\ell_{\text{attr}}(\bm{y},f_{A}(x;\theta_{A}))+\ell_{\text{cat}}(\bm{y},f_{A}(x;\theta_{A}))\end{split}$ (3)
@@ 4 Implementation and Experiments
We base our evaluation on three existing datasets: CLEVR (Johnson et al., 2017a), COCO (Lin et al., 2014), and GQA (Hudson and Manning, 2019), where their concepts are formed from original and supplemental annotations (Patterson and Hays, 2016). COCO and GQA contain both attribute and category concepts, while CLEVR provides only attribute concepts. For consistency across datasets, the shape attribute family in CLEVR is reinterpreted as category concepts when required.
In COCO and GQA, many annotated attributes appear extremely infrequently (sometimes only a few dozen instances), which provides insufficient training signal. To ensure meaningful supervision, we focus on the most frequent and well-represented concepts, retaining the top 35 attribute concepts for COCO and an equally sized set of 35 attribute concepts for GQA, together with their associated categories. This selection yields 64 total concepts for COCO and 68 for GQA.
To evaluate the scalability and generalization capability of our framework, we also construct a Unified dataset that aggregates all concepts and modality-specific representations from CLEVR, COCO, and GQA. This dataset contains overlapping concepts across domains, enabling us to examine how well the framework transfers and reconciles abstract concepts when they appear in different visual or linguistic forms. More details on the datasets and preprocessing steps can be found in Appendix C. Our experiments follow the same train and validation splits as the original datasets: the framework is pretrained on the train sets and evaluated on the validation sets.
@@ 4.1 Pretraining
Concept Space. To ensure that each concept box always has a valid set of lower boundaries smaller than its upper boundaries, we use two vectors, $(\omega_{\text{min},y},\omega_{\Delta,y})=\Omega_{y}$, instead of $(\omega_{\text{min}},\omega_{\text{max}})$ to represent a box in our actual experiments, where $\omega_{\Delta}\in\mathcal{K}_{\geq 0}$ is restricted to non-negative values. A box’s upper boundaries can be obtained by $\omega_{\text{max}}=\omega_{\text{min}}+\omega_{\Delta}$. We set the dimension of $\mathcal{K}$ to 50, based on empirical experiments (see Appendix A.5 for an ablation study on concept space dimensionality). Initial parameters for $\mathcal{C}$ are sampled from two uniform distributions. As for the negative sampling method, in CLEVR, the only negative concept pairs come from combinations of concepts residing in the same-attribute families, such as (red, blue) in the color family. For COCO and GQA, negative samples are randomly selected from all concepts. The concept space is trained for just two epochs for each dataset with a batch size of 256 using an AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $10^{-3}$. The training of this concept space can be completed quickly as there are only thousands of parameters for a moderately-sized concept space.
Projection Models. In adapting our framework to the datasets featuring vision and natural language modalities, we incorporate a vision projection model $f_{\text{vision}}$ based on a Vision Transformer encoder (Dosovitskiy et al., 2020) and a natural language projection model $f_{\text{NL}}$ based on a BERT encoder (Devlin et al., 2018). Both models utilize their encoders’ outputs on [CLS] tokens to generate projection boxes in $\mathcal{K}$. The outputs $e$ with a dimension of 768 are divided into two equal chunks, $h_{\text{min}}$ and $h_{\Delta}$, each with a dimension of 384. These chunks are then input into two fully connected layers to produce $\omega_{\text{min}}$ and $\omega_{\Delta}$ for their respective projection boxes. To ensure $\omega_{\Delta}$ is always a non-negative vector, an additional ReLU layer is applied. The complete projection process for inputs from the vision modality is outlined in Algorithm 1.
@@
[…]
@@
$h_{\text{min}}^{\text{vision}},h_{\Delta}^{\text{vision}}\leftarrow\text{split}(e_{\text{vision}})$
$\omega^{\text{vision}}_{\text{min}}\leftarrow\text{Linear}_{\text{min}}^{\text{vision}}(h_{\text{min}})$
@@
$\Omega^{\text{vision}}=(\omega^{\text{vision}}_{\text{min}},\omega^{\text{vision}}_{\Delta})$
@@
For each object $i$ in the CLEVR dataset, its attribute prediction for a specific attribute family $\bm{z}$ (e.g., color) is generated by $\bar{y}_{i}^{z}=\text{argmax}_{y\in\bm{z}}P(\Omega_{y}|\Omega_{i})$. For each object $i$ in COCO and GQA, a threshold is applied to $P(\Omega_{y}\mid\Omega_{i}),y\in\mathcal{Y}^{\text{attr}}$ to obtain attribute predictions, and category prediction is generated by $\bar{y}_{i}^{\text{cat}}=\text{argmax}_{y\in\mathcal{Y}^{\text{cat}}}P(\Omega_{y}\mid\Omega_{i})$.
@@
We establish a baseline by replacing the concept space with a traditional Multilayer Perceptron (MLP) at the classification head of $f_{\text{vision}}$. Additionally, we implement the vision-modality projection model using a ResNet model (He et al., 2015) as the backbone to showcase the flexibility of the proposed framework. Results summarized in Table 1 show that our proposed framework achieves comparable performance to traditional models while leveraging a novel concept space with interpretable learned knowledge.
Backbone Method CLEVR COCO GQA Unified
Acc Acc mAP Acc mAP Acc mAP
ResNet Baseline 0.997 0.894 0.532 0.725 0.202 0.897 0.447
$f_{\text{vision}}$ 0.990 0.892 0.520 0.727 0.191 0.892 0.439
ViT Baseline 0.999 0.960 0.594 0.842 0.335 0.949 0.512
$f_{\text{vision}}$ 0.999 0.957 0.590 0.844 0.352 0.949 0.502
Table 1: Comparison with baseline models on vision-modality classification. CLEVR uses Accuracy. COCO, GQA, and Unified evaluate category concepts using Accuracy and attribute concepts using mean Average Precision (mAP).
@@
Apart from featuring a concept-centric learning scheme, the proposed framework can also learn modality-specific knowledge faster by referencing learned knowledge from the modality-agnostic concept space as indicated in Fig. 2. This more natural learning process of our framework bridges the efficiency gap between traditional machine learning methods, which often demand extensive data, and human learning, which excels at adeptly and efficiently extracting modality-specific representations and associating them with mental entities of abstract knowledge. To fully evaluate the impact of this transparent, modality-agnostic concept space on the learning of modality-specific projection models, we conduct an ablation study on it in Sec. A.4.
@@
Figure 2: Learning curves of the proposed projection models and the baseline model. The plot displays the evaluation accuracy on category concepts and the evaluation mAP on attribute concepts measured every 50 training steps. During the learning process, the proposed vision-modality projection model improves more quickly compared to the baseline thanks to the universal concept space that already has abstract knowledge embedded in it. This faster learning process of our framework bridges the efficiency gap between traditional machine learning methods, which require a large amount of data, and human learning that excels at extracting modality-specific representations and linking them to structured abstract knowledge.
@@
Projection models for the natural-language modality achieve highly accurate performance ($\geq 99\%$) thanks to the clearly structured description sentences. Further implementation and training details of projection models can be found in Appendix D.
4.2 Cross Modality Alignment
A key distinction between our framework and most existing multimodal models is that each modality-specific projection model is trained independently, yet they remain inherently compatible and aligned because they share a common concept space that already encodes abstract knowledge. As a result, the projections from different modalities are well aligned without requiring any extensive joint training typically needed in models such as ViLT, CLIP, or FLAVA. We assess alignment between the vision and language modalities from our decoupled projection models by measuring the agreement between their knowledge space projections and by comparing the results with strong multimodal benchmark models. Because cross-modality alignment cannot be measured directly, we use accuracy on an image–text matching task (Section 4.4.1) from the Unified dataset with a large pool of 130 concepts as a proxy, and we initialize all comparison models with their pretrained weights to ensure fairness.
Before any fine-tuning, we measure each method’s initial alignment. Table 2 shows that our independently trained projection models already achieve strong cross-modal alignment at initialization, reaching over 95% accuracy without further training. We then apply a joint training objective from Appendix B to finetune the two projection models. This joint training is used only to obtain a direct comparison of cross-modality alignment efficiency against existing multimodal models and is not required for our framework to function. Figure 3 plots image–text matching accuracy every 50 steps against the GPU time elapsed during finetuning. Our framework not only starts with strong alignment but also converges significantly more quickly than existing multimodal models during joint training.
Model Init. Acc. Best Acc. Time to 95% Acc. (s)
CLIP 0.767 0.9701 42.7
FLAVA 0.499 0.9848 103.1
ViLT 0.747 0.9568 265.1
Ours 0.954 0.9568 – (already >>95% at step 1)
Table 2: Multimodal alignment comparison across vision-language models. We use image-text matching accuracy for validation split of the Unified dataset as a proxy for evaluating cross-modality alignment. Our framework achieves high alignment at initialization (over 95% accuracy at step 1) due to the shared knowledge, and requires significantly less computation than CLIP, FLAVA, or ViLT to reach their best performance.
020204040606080801001001201201401401601601801802002000.60.60.80.811Training GPU time (s)Valuation AccuracyOursViLTCLIPFLAVA
Figure 3: Comparison of training efficiency for cross-modality alignment. Validation accuracy for image-text matching task of the Unified dataset is used as a proxy for evaluating cross-modality alignment. Our proposed method reaches near-saturated alignment without finetuning and attains its final accuracy within a significantly shorter period, while ViLT, CLIP, and FLAVA require substantially more computation time to achieve comparable alignment.
4.3 Incorporating New Modality
Thanks to the decoupled training scheme of each individual modality-specific projection model and the inherent compatibility and alignment that result from sharing a unified knowledge space, incorporating new modalities into our framework is easy and scalable. To add a new modality to a functioning system with existing projection models, the new projection model only needs to be trained using the same projection-model training objective, and none of the existing modality-specific models need to be adjusted or jointly finetuned.
To demonstrate this scalable design, we translated the natural language descriptions originally used to train the BERT-based projection model into three new languages, including Chinese, French, and Spanish. Each translated language represents a new modality with a distinct input space added to a system that previously contained only vision and English (previously referred to as the natural language modality). Although all inputs are natural language, each language is treated as an independent modality because it is mapped into the shared knowledge space via its separately trained modality-specific projection model, rather than sharing parameters with the English text projection as in multilingual encoders. We follow the same training process as for the English projection model and train these new models independently while adapting them to the shared knowledge space. We then directly measure cross-modality alignment using a cross-modality representation-pair matching task made from the Unified dataset as a proxy. Results in Table 3 show strong alignment across all modalities. To gain further insight into the agreement between different projection models, Figure 4 reports the average entailment probabilities for all positive cross-modality representation pairs (two representations correspond to the same set of concepts) and negative ones (perturbed concepts) across all modality combinations.
This ability to incorporate new modalities without any joint finetuning with existing modalities highlights the flexibility and scalability of our framework. In addition, the modality-specific projection models for these new modalities demonstrate the same efficient training behavior as shown in Figure 2.
Modality Image English Chinese Spanish French
Image 0.955 0.968 0.961 0.935
English 0.955 0.974 0.982 0.876
Chinese 0.968 0.974 0.970 0.887
Spanish 0.961 0.982 0.970 0.878
French 0.935 0.876 0.887 0.878
Table 3: Cross-modality matching accuracy across image, English, Chinese, Spanish, and French. Each modality-specific model is trained independently with no joint training.
Figure 4: Modality alignment measured by entailment probabilities between positive and negative cross-modality representation pairs across image, English, Chinese, Spanish, and French modalities. Left: higher values (↑) indicate stronger projection overlap for positive concept pairs. Right: lower values (↓) indicate stronger separation for negative pairs. All modality-specific models are trained independently.
4.4 Downstream Tasks